OpenStudy (jenniferjuice):

Identify the correct slope and y-intercept of the equation y - 4x = -3

OpenStudy (jenniferjuice):

slope = -1; y-intercept = (0, -3) slope = -1/4; y-intercept = (0, 3) slope = 4; y-intercept = (0, -3) slope = 4; y-intercept = (0, 3)

OpenStudy (anonymous):

Do you know slope intercept form?

OpenStudy (anonymous):

you need to put your equation in y = mx + b form. Can you do that ?

OpenStudy (anonymous):

Once you do that, your slope will be in the m spot and your y intercept will be in the b spot :)

OpenStudy (anonymous):

Kellie is right

OpenStudy (anonymous):

You can do this Jennifer...it is not that hard

OpenStudy (anonymous):

Come on Jennifer. You can. We know you can. YOU CAN DO IT

OpenStudy (anonymous):

If melody and me can do this, so can you

OpenStudy (anonymous):

Yes, many people have done it before. You will be part of that group. Have Faith!

OpenStudy (anonymous):

just remember with your y intercept, make sure it is in the y spot (x,y)

OpenStudy (anonymous):

remember the equation is y= mx +b

OpenStudy (anonymous):

oh well melody...we did what we could without giving the answer

OpenStudy (anonymous):

isolate the Y by undoing the operation

OpenStudy (anonymous):

Yes Kellie, we did good.

OpenStudy (anonymous):

she will get to it when she can...and if she needs help she can tag us :)

OpenStudy (anonymous):

Egg- zack-ly

OpenStudy (anonymous):

lol

OpenStudy (jenniferjuice):

you guys are a great team

OpenStudy (jenniferjuice):

lol thanks i got it ;)

OpenStudy (jenniferjuice):

*:)

OpenStudy (anonymous):

Happy to help :)

OpenStudy (anonymous):

just curious...what answer did you get ?